Solve by Elimination or substitution Method

3x - 2y = 8
-12x + 8y = 32

Answer by oscargut(2103)

3x - 2y = 8 (x4)
-12x + 8y = 32
12x-8y=32
-12x + 8y = 32
adding both eq
0=64 then the system is inconsistent (no solution)